EVA Report – January 10th

Author: Cesare Guariniello

Purpose of EVA: Testing the next phase of the Yagi-Uda antenna-based navigation experiment for navigating back to the hab. Testing with a stranded astronaut in a condition of low visibility

Destination: End of Mountain Goat Road

UDM27 Coordinates:

519750 E

4250500 N

Participants: Cesare, Melanie, Mark (stranded astronaut)

Narrative: This EVA put an astronaut in condition of low visibility, simulating a dust storm. This condition was obtained with a cardboard placed over the top portion of the ExoSuit helmet. The astronaut has visibility limited to about 5m ahead (for reasons of safety) and can look at the handheld radio when navigating, in order to find the direction of the habitat. The crew egressed the habitat with Mark already “under the hood”, and checked the safety of the hooding apparatus, then the three EVA crewmembers started walking Northwest to the intended destination, guided by Cesare’s GPS unit. During the outward path, Melanie and Cesare had Mark stopping multiple times, to spin around and be redirected so as to be completely disoriented. The outward path was also chosen so that the return direction would not follow it.

Once the turnaround point was reached, Mark turned on the Yagi-Uda antenna receiver and began scanning for the radio beacon from the habitat, sent initially at intervals of 5 minutes. Melanie and Cesare took care of the safety of the path chosen by Mark, but let him decide the direction to walk towards. When rock walls occurred, Mark decided the best path around them, and only once he was requested to stop to avoid excessive proximity to a shallow canyon. The EVA gave significant results to the experiment. Mark was able to locate the habitat with adequate precision even when hills and rocks obstructed the line of communication, though a power attenuator might be required in close proximity to the habitat (Mark could navigate close to saturation of the signal, at about 500m from the habitat). A major problem was the difficulty to follow a straight path. Mark always pinpointed the direction towards the habitat, but walked in wide arcs. For this reason, Cesare and Melanie requested continuous transmission of the beacon signal which helped reducing the path error.

Around 1:40 (one hour and 35 minutes into the EVA), the weather turned to rain. The crew immediately reported the occurrence to the HabCom. Upon indication by the commander the EVA was scrubbed immediately, Mark was unhooded, and the EVA team returned to the hab (which had been in sight for about 20 minutes, and was only 500m away) without delays and ingresses the airlock after collecting two soil samples from the immediate vicinity of the habitat, to be analyzed for use in the Greenhab by future crews.

Cesare Guariniello, crew geologist

Crew 186 – Boilers2Mars

Mars Desert Research Station

Operations Report – January 10th

Crew

186 Operations Report 10JAN2018

SOL:

10

Name

of person filing report: M. Grande

Non-nominal

systems: None

Notes

on non-nominal systems: None

Generator

(hours run): 12.7h

Generator

batteries switched at 8:45am

Generator

batteries switched at 7:30pm

Solar—

SOC

@

7:15am : 95%

@

6:00pm : 99%

Diesel:

50%

Propane:

29%

Ethanol

Free Gasoline (5 Gallon containers for ATV): 5.9 Gallons

Water

(trailer): 10 Gallons

Water

(static): 413 Gallons

Trailer

to Static Pump used: No

Water

(loft) – Static to Loft Pump used: Yes

Water

Meter: 129321.3 Gallons

Toilet

tank emptied: Yes

ATVs

Used: None

Oil

Added? No

ATV

Fuel Used: 00 Gallons

#

Hours the ATVs were used today: 00:00 hours

Notes

on ATVs: None.

Deimos

rover used: No

Hours:

99.0

Beginning

charge: 100%

Ending

charge:

Currently

charging: Yes

Sojourner

rover used: Assigned to director only.

Hours:

Director discretional hours

Beginning

charge: 100%

Ending

charge:

Currently

charging: Yes

Spirit

rover used: No

Hours:

10.5

Beginning

charge: 95%

Ending

charge:

Currently

charging: Yes

Opportunity

rover used: No

Hours:

5.5

Beginning

charge: 100%

Ending

charge:

Currently

charging: No

Curiosity

rover used: No

Hours:

3.1

Beginning

charge: 100%

Ending

charge:

Currently

charging: No

HabCar

used and why, where? No

General

notes and comments: None.

Summary

of internet: All nominal

Summary

of suits and radios: All nominal

Summary

of Hab operations: All nominal

Summary

of GreenHab operations: All nominal

Summary

of ScienceDome operations: All nominal

Summary

of RAM operations: Not Operational

Summary

of health and safety issues: Crew is Healthy

Questions,

concerns and requests to Mission Support: I will wait for the handover day to review the state of the generator coolant as well as some other things, like the oil and filters. Curious when the last time the generator was completely serviced? (i.e. air and

oil and fuel filters replaced, hoses and tubing checked, etc.) I believe the man who installed or managed the settings for our power system will be coming back to town soon and will hopefully address the concerns we have with the system, and I would suggest

that a generator servicing might be scheduled around the same time, if that hasn’t been done recently.

As mentioned before: we need a propane resupply, the right tailgate latch on Opportunity is broken, and there is an unidentified leak underneath the front of the hab.

Regards,

Melanie

Grande, Crew Engineer, Crew 186

Greenhab Report – January 10th

GreenHab Report

Mark Gee

10Jan2018

Environmental control:

Heating

Shade cloth on

Working Hour: 06:50PM

Inside temp at working hour: 18 C

Outside temp during working hours: 0 C

Inside temperature high: 26 C

Inside temperature low: 15 C

Inside humidity: 46 %RH

Inside humidity high: 60 %RH

Inside humidity low: 34 %RH

Hours of supplemental light:

For the crops 05:00 to 11:59 PM

Changes to crops: Microgreens almost ready to harvest. Spinach is growing rapidly. Cucumbers continue to produce fruit. Beans near harvest. Tomatoes are flowering but no fruit set. Sprouts are emerging from the seeds planted a few days ago.

Daily water usage for crops: 10 gallons

Time(s) of watering for crops: 06:30PM

Research observations: Microgreens are growing rapidly. No change to tomatoes sprayed with moringa extract.

Changes to research plants: None.

Aquaponics: Not in use.

Narrative: Growing up, we learn to fear bees because of their sting, then to like them because of their honey. But we should treasure them for their work as pollinators. Bees and other insects perform the critical task of transferring pollen from flower to flower. Without this transfer by insect, many crops would not produce fruit because their pollen will not blow in the wind. Some crops, namely the Cucurbitacea family of pumpkins, melons, squash and cucumbers, have co-evolved with their own species of bees that specialize in pollinating these crops. Almond farmers will actually pay beekeepers to bring their bees into the almond orchard while the trees are flowering to help ensure that there will be a bountiful harvest.

Unfortunately, there are no bees on Mars.

To make up for this, I put on my bee suit in the Green Hab today, plucked a male flower from a cucumber plant, and buzzed from cucumber to cucumber, sprinkling pollen in the flower as I went. It took effort and patience.

There are many natural resources that we do not appreciate until they are gone. Surviving on a new planet will be challenging and full of surprising realizations of how much we are missing back on Earth.

Support/supplies needed: I have several questions of general interest.

Is the water supply from Hanksville drawn out of a well, river, or something else?

What are the heater control settings? What temperature is it set to turn on and what temperature for off?

What is to be done with used potting soil?

Is the shade cloth a 60/40 cloth? Is there a better way to describe how much shade it gives?

Science Report – January 10th

Science Report – Microbiology

10JAN2018

Author: Samuel Albert, Crew 186 Health & Safety Officer

Of the four DNA sequencing runs originally planned, three have been completed so far. The first run encountered errors and yielded poor results, only about 350 reads. The second run, which sampled from crops growing in the GreenHab, yielded much better results, over 600,000 reads. The third run, which sampled from locations on the upper deck of the habitat, yielded strong results as well, about 26,000 results. The fourth run, which will be completed in the last few days of the mission, is planned to sample from the bathroom and shower area on the lower deck of the habitat. Following the mission, all results will be analyzed to assess which microbes were found in the various sampling locations.

Journalist Report – January 10th

[Sol 10] [The
Lost Astronaut]

The day began with our usual order of activities – yoga and breakfast followed by EVA prep. The purpose of this EVA was a dedicated test of the radio navigation experiment. But unlike the previous tests, the astronaut using the antenna would have their vision restricted to only their immediate area by a cardboard visor strapped to their helmet. This, combined with the flat lighting of the gloomy sky, meant there was no possibility of using visual cues to return to the hab. Furthermore, the antenna would not be in the hands of her designer, but the crew’s greenhab scientist, Mark Gee, who possessed no previous experience using the antenna. It was out closest simulation yet of a real lost astronaut imperiled by low-visibility conditions.

Due to the risk of precipitation in the morning, the EVA team did not depart until slightly past noon when the weather began to stabilize. Cesare Guariniello and Melanie Grande supported Mark on EVA (i.e. ensured he wouldn’t accidentally walk off a cliff) and led him on foot to the east until he was thoroughly disoriented. At 1 pm, we switched on the habitat’s navigation beacon. Thirty minutes later, the EVA team emerged from behind a distant ridge within sight of the habitat. Melanie and Cesare later reported that they struggled to keep up with Mark as he aggressively chased the signal back to the hab. But before the test could proceed to completion, a freezing Martian snow began trickling from the skies. By mission rule the team was forced to abort and return immediately.

The team made use of the extra time to relax and prepare for an evening teleconference with the Purdue chapter of the Mars Society. It went splendidly. Since then, Mark has been cross pollinating plants in the greenhab, and Sam Albert, the health and safety officer, made preparations to take additional microbial samples tomorrow. The rest of us are gearing up for another test of the navigation antenna soon.

Justin Mansell, MDRS Crew 186 Journalist

P.S. Photos attached. Photo of the day: 10Jan2018 Preparing the lost astronaut.jpg